OpenStudy (anonymous):

Electricity goes about $0.05 Kilowatt hr, in form of electricity how many joules can you get for a dollar?

OpenStudy (turingtest):

a joule is a watt times a second, so we have\[$1\times\frac{10^3W\text{hr}}{$0.05}\times\frac{3600s}{\text{hr}}\]will give us watt-seconds=joules

OpenStudy (anonymous):

So my answer is 7.2 x 10^7 J is that right?

OpenStudy (turingtest):

seems that way

OpenStudy (anonymous):

Thanks!